# MAPSS: Manifold-based Assessment of Perceptual Source Separation

### v1

@@ 3.2 Constructing Perceptual Clusters on the Manifold
Let us recall $N^{l}_{f}$ from Section 2. Given $i\in\{0,\dots,N^{l}_{f}-1\}$, the waveform reference signal of the $i$-th source, $\mathbf{y}_{i}$, undergoes $N_{p}$ hand-crafted perceptual distortions detailed in Appendix D. We define the $i$-th distortion set $\mathcal{D}_{i}$ as:
$\mathcal{D}_{i}=\big{\{}\mathbf{\hat{y}}_{i},\mathbf{y}_{i},\mathbf{y}_{i,1},\ldots,\mathbf{y}_{i,{N_{p}}}\big{\}},$ (17)
where $\mathbf{y}_{i,p}\in\mathbb{R}^{L}$ is the $p$-th perceptual distortion of $\mathbf{y}_{i}$, for $p\in\{1,\ldots,N_{p}\}$}. Each waveform in $\mathcal{D}_{i}$ is independently encoded via a pre-trained self-supervised model, e.g., wav2vec 2.0 Baevski et al. (2020). Let $\Phi:\mathbb{R}^{L}\to\mathbb{R}^{M}$ be this encoding operator, then:
$\mathbf{x}_{i,p}=\Phi\left(\mathbf{y}_{i,p}\right),\quad\mathbf{x}_{i}=\Phi\left(\mathbf{y}_{i}\right),\quad\hat{\mathbf{x}}_{i}=\Phi\left(\hat{\mathbf{y}}_{i}\right).$ (18)
Applying (17), (18) across all $N^{l}_{f}$ speech sources, we get the complete high-dimensional dataset:
$\mathcal{X}=\Big{\{}\hat{\mathbf{x}}_{i},\mathbf{x}_{i},\mathbf{x}_{i,1},\dots,\mathbf{x}_{i,{N_{p}}}\;\Big{|}\;i=0,\dots,N^{l}_{f}-1\Big{\}},$ (19)
with $|\mathcal{X}|=N^{l}_{f}(N_{p}+2):=N$. The diffusion maps embedding of $\mathbf{x}\in\mathcal{X}$ on the manifold is thus:
$\Psi_{t}^{(d)}(\mathbf{x})=\left(\lambda_{1}^{t}\mathbf{u}_{1}(k),\dots,\lambda_{d}^{t}\mathbf{u}_{d}(k)\right)^{T}\in\mathbb{R}^{d},$ (20)
where $k\in\big{\{}1,\ldots,N\big{\}}$ is the global index of $\mathbf{x}$. We can now define the $i$-th perceptual cluster on the manifold in the truncated dimension $d$ as:
$c_{i}^{(d)}=\left\{\Psi_{t}^{(d)}(\mathbf{x}_{i}),\Psi_{t}^{(d)}(\mathbf{x}_{i,p})\;\middle|\;p=1,\dots,N_{p}\right\}.$ (21)
where we exclude the system output embedding $\Psi_{t}^{(d)}(\mathbf{\hat{x}}_{i})$ from $c_{i}^{(d)}$, since this embedding will be measured against the cluster statistics to yield the PS and PM measures. Including it in the cluster statistics would create circular dependency, where the embedding being evaluated would influence its own evaluation metrics.
By creating a group of waveform distortions from each reference, and then embedding their encoded versions using diffusion maps, perceptual geometric clusters are formed around the references on the manifold, as demonstrated in Figure 1. Crucially, Euclidean distances between each pair of embedded coordinates align with the diffusion distance between their corresponding self-supervised representations (14). We hypothesize and empirically demonstrate that these geometric distances can also be backtracked to perceptual dissimilarity between the corresponding waveforms. This perceptual-geometric correspondence forms the theoretical foundation for our proposed PS and PM measures.
Figure 1: Two embedded clusters on the manifold, associated with a two-speaker English mixture, created via diffusion maps. Each cluster consists of the reference of its speaker denoted by ‘x’ and its distortions marked by circles. We use the two prime coordinates $\left(\lambda_{1}^{t}\mathbf{u}_{1},\,\lambda_{2}^{t}\mathbf{u}_{2}\right)$ with $t=1$ (15).
@@ 3.3 The Perceptual Separation (PS) Measure
We aim to quantify the perceptual separation of $\mathbf{\hat{y}}_{i}$ from its non-attributed references $\{\mathbf{y}_{j}\}_{i\neq j}$ with the Mahalanobis distance Evans et al. (2021). The empirical centroid and unbiased covariance matrix of the cluster $c^{(d)}_{j}$ are:
$\hat{\boldsymbol{\mu}}_{j}^{(d)}$ $=\frac{1}{\Big{|}c^{(d)}_{j}\Big{|}}\sum_{\boldsymbol{\psi}\in c_{j}^{(d)}}\boldsymbol{\psi},$ (22)
$\widehat{\boldsymbol{\Sigma}}_{j}^{(d)}$ $=\frac{1}{\Big{|}c^{(d)}_{j}\Big{|}-1}\sum_{\boldsymbol{\psi}\in c_{j}^{(d)}}\left(\boldsymbol{\psi}-\hat{\boldsymbol{\mu}}_{j}^{(d)}\right)\left(\boldsymbol{\psi}-\hat{\boldsymbol{\mu}}_{j}^{(d)}\right)^{T},$ (23)
@@
where $\hat{\boldsymbol{\mu}}_{j}^{(d)}\in\mathbb{R}^{d},\text{ }\widehat{\boldsymbol{\Sigma}}_{j}^{(d)}\in\mathbb{R}^{d\times d}$. The Mahalanobis distance from the embedding of the $i$-th output $\boldsymbol{\Psi}_{t}^{(d)}(\hat{\mathbf{x}}_{i})$ to $c_{j}^{(d)}$ is given by:
$d_{M}\left(\boldsymbol{\Psi}_{t}^{(d)}(\hat{\mathbf{x}}_{i});\hat{\boldsymbol{\mu}}_{j}^{(d)},\widehat{\boldsymbol{\Sigma}}_{j}^{(d)}\right)=$ (24)
$\sqrt{\left(\boldsymbol{\Psi}_{t}^{(d)}(\hat{\mathbf{x}}_{i})-\hat{\boldsymbol{\mu}}_{j}^{(d)}\right)^{T}\left(\widehat{\boldsymbol{\Sigma}}_{j}^{(d)}+\epsilon I^{(d)}\right)^{-1}\left(\boldsymbol{\Psi}_{t}^{(d)}(\hat{\mathbf{x}}_{i})-\hat{\boldsymbol{\mu}}_{j}^{(d)}\right)},$
@@
where we use for regularization $\epsilon=10^{-6}$ with the $d$-dimensional identity matrix $I^{(d)}$. We define the measured Mahalanobis distance from $\boldsymbol{\Psi}_{t}^{(d)}(\hat{\mathbf{x}}_{i})$ to its attributed and closest non-attributed clusters as:
$\hat{A}_{i}^{(d)}$ $=d_{M}\left(\boldsymbol{\Psi}_{t}^{(d)}(\hat{\mathbf{x}}_{i});\hat{\boldsymbol{\mu}}_{i}^{(d)},\widehat{\boldsymbol{\Sigma}}_{i}^{(d)}\right),$ (25)
$\hat{B}_{i}^{(d)}$ $=\min_{j\neq i}d_{M}\left(\boldsymbol{\Psi}_{t}^{(d)}(\hat{\mathbf{x}}_{i});\hat{\boldsymbol{\mu}}_{j}^{(d)},\widehat{\boldsymbol{\Sigma}}_{j}^{(d)}\right),$ (26)
Then, the measured PS score for $\mathbf{\hat{y}}_{i}$ in the truncated dimension $d$ is:
$\widehat{\text{PS}}_{i}^{(d)}=1-\frac{\hat{A}_{i}^{(d)}}{\hat{A}_{i}^{(d)}+\hat{B}_{i}^{(d)}},$ (27)
where higher values are desirable and by design $\hat{A}_{i}^{(d)}+\hat{B}_{i}^{(d)}>0$. The distance to the attributed cluster $\hat{A}_{i}^{(d)}$ and to the nearest foreign cluster $\hat{B}_{i}^{(d)}$ balance fidelity and separation, as demonstrated in Figure 2. The PS maintains high values whenever $\hat{A}_{i}^{(d)}\ll\hat{B}_{i}^{(d)}$, and falls as $\hat{A}_{i}^{(d)}$ outgrows $\hat{B}_{i}^{(d)}$, representing growing leakage when the output of the $i$-th source in the separation system may perceptually resemble other references more than its attributed one.
Since within each cluster the distribution of distortions may not be isotropic, i.e., some directions capture subtle perturbations while others represent more variable changes, then using Euclidean distance alone would ignore this anisotropy and will not emphasize distortions depending on direction. The Mahalanobis distance resolves this by normalizing with the empirical covariance of each cluster, weighting deviations more heavily in stable directions and less in noisy ones. In this sense, Mahalanobis does not contradict the property (14), it refines it locally and preserves it by incorporating second-order structure of the manifold.
@@ 3.4 The Perceptual Match (PM) Measure
The PM measure aims to quantify how perceptually aligned the estimated output $\hat{\mathbf{y}}_{i}$ is with its reference $\mathbf{y}_{i}$. Let $\tilde{c}_{i}^{(d)}=c_{i}^{(d)}\setminus\Psi_{t}^{(d)}(\mathbf{x}_{i})$ denote the reference-free $i$-th cluster. We compute the unbiased empirical covariance matrix of $\tilde{c}_{i}^{(d)}$, relative to its reference embedding:
$\widehat{\widetilde{\boldsymbol{\Sigma}}}_{i}^{(d)}=\frac{1}{\big{|}\tilde{c}^{(d)}_{i}\big{|}-1}\sum_{\mathbf{\boldsymbol{\psi}}\in\tilde{c}_{i}^{(d)}}\left(\boldsymbol{\psi}-\boldsymbol{\Psi}_{t}^{(d)}(\mathbf{x}_{i})\right)\left(\boldsymbol{\psi}-\boldsymbol{\Psi}_{t}^{(d)}(\mathbf{x}_{i})\right)^{T}.$ (28)
Then, for $p\in\{1,\ldots,N_{p}\}$ the squared Mahalanobis distance from the $p$-th distortion to its attributed reference in the $i$-th cluster, is given by:
$d^{2}_{M}\left(\boldsymbol{\Psi}_{t}^{(d)}(\mathbf{x}_{i,p});\boldsymbol{\Psi}_{t}^{(d)}(\mathbf{x}_{i}),\widehat{\widetilde{\boldsymbol{\Sigma}}}_{i}^{(d)}\right)=$ (29)
$\hskip-60.00009pt\left(\boldsymbol{\Psi}_{t}^{(d)}(\mathbf{x}_{i,p})-\boldsymbol{\Psi}_{t}^{(d)}(\mathbf{x}_{i})\right)^{T}\left(\widehat{\widetilde{\boldsymbol{\Sigma}}}_{i}^{(d)}+\epsilon I^{(d)}\right)^{-1}\left(\boldsymbol{\Psi}_{t}^{(d)}(\mathbf{x}_{i,p})-\boldsymbol{\Psi}_{t}^{(d)}(\mathbf{x}_{i})\right),$
where as in the PS case, $\epsilon=10^{-6}$. Let us define the set of distances:
$\hat{\mathcal{G}}^{(d)}_{i}=\left\{d^{2}_{M}\left(\boldsymbol{\Psi}_{t}^{(d)}(\mathbf{x}_{i,p});\boldsymbol{\Psi}_{t}^{(d)}(\mathbf{x}_{i}),\widehat{\widetilde{\boldsymbol{\Sigma}}}_{i}^{(d)}\right)\;\middle|\;p=1,\dots,N_{p}\right\}.$ (30)
Empirically, we observed that nearly always these distances are well-approximated by a Gamma distribution, validated using Kolmogorov-Smirnov goodness-of-fit tests Kolmogorov (1986); Smirnov (1948). The sample mean and unbiased variance of $\hat{\mathcal{G}}^{(d)}_{i}$ are estimated by:
[…]
$\hat{\sigma}^{2}_{\mathcal{G}^{(d)}_{i}}=\frac{1}{\Big{|}\hat{\mathcal{G}}^{(d)}_{i}\Big{|}-1}\sum_{g\in\hat{\mathcal{G}}^{(d)}_{i}}{\left(g-\hat{\mu}_{\mathcal{G}^{(d)}_{i}}\right)^{2}},$ (32)
@@
$\hat{k}^{(d)}_{i}=\frac{\hat{\mu}_{\mathcal{G}^{(d)}_{i}}^{2}}{\hat{\sigma}^{2}_{\mathcal{G}^{(d)}_{i}}},$ (33)
$\hat{\theta}^{(d)}_{i}=\frac{\hat{\sigma}^{2}_{\mathcal{G}^{(d)}_{i}}}{\hat{\mu}_{\mathcal{G}^{(d)}_{i}}}$. (34)
Similarly, the squared Mahalanobis distance from the output embedding to its attributed cluster is:
$d^{2}_{M}\left(\boldsymbol{\Psi}_{t}^{(d)}(\hat{\mathbf{x}}_{i});\boldsymbol{\Psi}_{t}^{(d)}(\mathbf{x}_{i}),\widehat{\widetilde{\boldsymbol{\Sigma}}}_{i}^{(d)}\right):=\hat{a}_{i}^{(d)}.$ (35)
@@
Consider $Q(k,x)=\Gamma(k,x)/\Gamma(k)$ as the regularized upper incomplete Gamma function. Then, the PM score for $\hat{\mathbf{y}}_{i}$ in dimension $d$ is:
$\widehat{\text{PM}}^{(d)}_{i}=Q\!\left(\hat{k}^{(d)}_{i},\frac{\hat{a}_{i}^{(d)}}{\hat{\theta}^{(d)}_{i}}\right)$, (36)
where again a higher score is more desirable, and $\hat{k}^{(d)}_{i}$ and $\hat{\theta}^{(d)}_{i}$ are well-defined by design as long as $N_{p}\geq 1$.

### v2

@@ 4.1 Constructing Perceptual Clusters on the Manifold
The waveform reference signal of the $i$-th source, $\mathbf{y}_{i}$, undergoes $N_{p}$ perceptual distortions, e.g., noise gating in different thresholds, vibrato in various rates, and a comb filter with several delay-gain pairs. Typically, $N_{p}\in\left[60,70\right]$. We define the $i$-th distortion set $\mathcal{D}_{i}$ as:
$\mathcal{D}_{i}=\big\{\mathbf{\hat{y}}_{i},\mathbf{y}_{i},\mathbf{y}_{i,1},\ldots,\mathbf{y}_{i,{N_{p}}}\big\},\quad\forall p\in\{1,\ldots,N_{p}\}:\mathbf{y}_{i,p}\in\mathbb{R}^{L},$ (12)
with $L$ from (1). Each waveform in $\mathcal{D}_{i}$ is independently encoded via a pre-trained self-supervised model, e.g., wav2vec 2.0 (Baevski et al., 2020). Let $\Phi:\mathbb{R}^{L}\to\mathbb{R}^{M}$ be this encoding operator, with $M$ from §3, so $\mathbf{x}_{i,p}=\Phi\left(\mathbf{y}_{i,p}\right),\,\mathbf{x}_{i}=\Phi\left(\mathbf{y}_{i}\right),\,\hat{\mathbf{x}}_{i}=\Phi\left(\hat{\mathbf{y}}_{i}\right)$}. Applying (12) across all $N_{f}$ sources results in the high-dimensional set of representations:
$\mathcal{X}=\Big\{\hat{\mathbf{x}}_{i},\mathbf{x}_{i},\mathbf{x}_{i,1},\dots,\mathbf{x}_{i,{N_{p}}}\;\Big|\;i=1,\dots,N_{f}\Big\},$ (13)
with $|\mathcal{X}|=N_{f}\,(N_{p}+2):=N$. We define the $i$-th perceptual cluster $\mathcal{C}_{i}^{(d)}$ on manifold $\mathcal{M}^{\left(d\right)}$ (10):
$\mathcal{C}_{i}^{(d)}=\left\{\bm{\Psi}_{t}^{(d)}(\mathbf{x}_{i}),\bm{\Psi}_{t}^{(d)}(\mathbf{x}_{i,p})\;\middle|\;p=1,\dots,N_{p}\right\}.$ (14)
where we exclude the embedding of the system output $\bm{\Psi}_{t}^{(d)}(\mathbf{\hat{x}}_{i})\in\mathbb{R}^{d}$ (10) from $\mathcal{C}_{i}^{(d)}$, since this embedding will be measured against the cluster statistics to produce the PS and PM measures. Including $\bm{\Psi}_{t}^{(d)}(\mathbf{\hat{x}}_{i})$ in the cluster would create a circular dependency that will bias the PS and PM. These distortions were hand-crafted to create a wide perceptual auditory coverage relative to the reference, e.g., by considering mildly-intrusive additive colored noise with signal-to-noise-ratios (SNRs) of 15 dB on one hand, and severely degrading heavy-tailed reverberations on the other hand.
[…]
@@ 4.2 The Perceptual Separation (PS) Measure
For source $i$, we aim to quantify the perceptual separation of $\mathbf{\hat{y}}_{i}$ from its non-attributed references $\{\mathbf{y}_{j}\}_{i\neq j}$ with the Mahalanobis distance (Evans et al., 2021). The empirical centroid and unbiased covariance matrix of the cluster $\mathcal{C}^{(d)}_{j}$ are:
$\hat{\bm{\mu}}_{j}^{(d)}=\frac{1}{\Big|\mathcal{C}^{(d)}_{j}\Big|}\sum_{\bm{\psi}\in\mathcal{C}_{j}^{(d)}}\bm{\psi},\quad\widehat{\bm{\Sigma}}_{j}^{(d)}=\frac{1}{\Big|\mathcal{C}^{(d)}_{j}\Big|-1}\sum_{\bm{\psi}\in\mathcal{C}_{j}^{(d)}}\left(\bm{\psi}-\hat{\bm{\mu}}_{j}^{(d)}\right)\left(\bm{\psi}-\hat{\bm{\mu}}_{j}^{(d)}\right)^{T},$ (15)
@@
where $\hat{\bm{\mu}}_{j}^{(d)}\in\mathbb{R}^{d},\text{ }\widehat{\bm{\Sigma}}_{j}^{(d)}\in\mathbb{R}^{d\times d}$. The squared Mahalanobis distance from the embedding of the $i$-th output $\bm{\Psi}_{t}^{(d)}(\hat{\mathbf{x}}_{i})$ to $\mathcal{C}_{j}^{(d)}$ is given by:
$d^{2}_{M}\left(\bm{\Psi}_{t}^{(d)}(\hat{\mathbf{x}}_{i});\hat{\bm{\mu}}_{j}^{(d)},\widehat{\bm{\Sigma}}_{j}^{(d)}\right)=\left(\bm{\Psi}_{t}^{(d)}(\hat{\mathbf{x}}_{i})-\hat{\bm{\mu}}_{j}^{(d)}\right)^{T}\left(\widehat{\bm{\Sigma}}_{j}^{(d)}+\epsilon I^{(d)}\right)^{-1}\left(\bm{\Psi}_{t}^{(d)}(\hat{\mathbf{x}}_{i})-\hat{\bm{\mu}}_{j}^{(d)}\right),$ (16)
@@
where we use for regularization $\epsilon=10^{-6}$ with the $d$-dimensional identity matrix $I^{(d)}$. We define the measured Mahalanobis distance from $\bm{\Psi}_{t}^{(d)}(\hat{\mathbf{x}}_{i})$ to its attributed and closest non-attributed clusters as:
$\hat{A}_{i}^{(d)}=d_{M}\left(\bm{\Psi}_{t}^{(d)}(\hat{\mathbf{x}}_{i});\hat{\bm{\mu}}_{i}^{(d)},\widehat{\bm{\Sigma}}_{i}^{(d)}\right),\quad\hat{B}_{i}^{(d)}=d_{M}\left(\bm{\Psi}_{t}^{(d)}(\hat{\mathbf{x}}_{i});\hat{\bm{\mu}}_{j^{*}}^{(d)},\widehat{\bm{\Sigma}}_{j^{*}}^{(d)}\right),$ (17)
with $j^{*}=\operatorname*{arg\,min}_{j\in\{1,\ldots,N_{f}\},\text{ }j\neq i}d_{M}\left(\bm{\Psi}^{\left(d\right)}_{t}\left(\hat{\mathbf{x}}_{i}\right);\bm{\mu}^{(d)}_{j},\bm{\Sigma}^{(d)}_{j}\right)$. Notice that (17) resembles the source permutation minimization processing in source separation evaluations (Le Roux et al., 2019).
The measured PS score for $\mathbf{\hat{y}}_{i}$ in the truncated dimension $d$ is:
$\widehat{\text{PS}}_{i}^{(d)}=1-\frac{\hat{A}_{i}^{(d)}}{\hat{A}_{i}^{(d)}+\hat{B}_{i}^{(d)}},\quad\widehat{\text{PS}}_{i}^{(d)}\in\left[0,1\right].$ (18)
where by design $\hat{A}_{i}^{(d)}+\hat{B}_{i}^{(d)}>0$ and a higher score is better. Functionally, when $\hat{A}_{i}^{(d)}\ll\hat{B}_{i}^{(d)}$ then the $i$-th output perceptually resembles its cluster members significantly more than competing cluster members and $\widehat{\text{PS}}_{i}^{(d)}$ approaches 1. $\hat{B}_{i}^{(d)}\ll\hat{A}_{i}^{(d)}$ indicates the opposite, and $\widehat{\text{PS}}_{i}^{(d)}$ drops towards 0.
@@ 4.3 The Perceptual Match (PM) Measure
The PM measure aims to quantify how perceptually aligned the estimated output $\hat{\mathbf{y}}_{i}$ is with its reference $\mathbf{y}_{i}$. Let $\tilde{\mathcal{C}}_{i}^{(d)}={\mathcal{C}}_{i}^{(d)}\setminus\Psi_{t}^{(d)}(\mathbf{x}_{i})$ denote the reference-free $i$-th cluster. Unlike Equation (15), we compute the unbiased empirical covariance matrix of $\tilde{\mathcal{C}}_{i}^{(d)}$ relative to its reference embedding:
$\widehat{\widetilde{\bm{\Sigma}}}_{i}^{(d)}=\frac{1}{\big|\tilde{\mathcal{C}}^{(d)}_{i}\big|-1}\sum_{\mathbf{\bm{\psi}}\in\tilde{\mathcal{C}}_{i}^{(d)}}\left(\bm{\psi}-\bm{\Psi}_{t}^{(d)}(\mathbf{x}_{i})\right)\left(\bm{\psi}-\bm{\Psi}_{t}^{(d)}(\mathbf{x}_{i})\right)^{T}.$ (19)
Then, for $p\in\{1,\ldots,N_{p}\}$, the squared Mahalanobis distance from the $p$-th distortion to its attributed reference in the $i$-th cluster, is given by $d^{2}_{M}\left(\bm{\Psi}_{t}^{(d)}(\mathbf{x}_{i,p});\bm{\Psi}_{t}^{(d)}(\mathbf{x}_{i}),\widehat{\widetilde{\bm{\Sigma}}}_{i}^{(d)}\right)$, following the definition in Equation (16). Let us define the set of distances:
$\hat{\mathcal{G}}^{(d)}_{i}=\left\{d^{2}_{M}\left(\bm{\Psi}_{t}^{(d)}(\mathbf{x}_{i,p});\bm{\Psi}_{t}^{(d)}(\mathbf{x}_{i}),\widehat{\widetilde{\bm{\Sigma}}}_{i}^{(d)}\right)\;\middle|\;p=1,\dots,N_{p}\right\}.$ (20)
Empirically, we validated that nearly always these distances are well-approximated by a Gamma distribution, using Kolmogorov-Smirnov goodness-of-fit tests (Smirnov, 1948; Kolmogorov, 1992). The sample mean and unbiased variance of $\hat{\mathcal{G}}^{(d)}_{i}$ are estimated by:
$\hat{\mu}_{\mathcal{G}^{(d)}_{i}}=\frac{1}{\Big|\hat{\mathcal{G}}^{(d)}_{i}\Big|}\sum_{g\in\hat{\mathcal{G}}^{(d)}_{i}}g,\quad\hat{\sigma}^{2}_{\mathcal{G}^{(d)}_{i}}=\frac{1}{\Big|\hat{\mathcal{G}}^{(d)}_{i}\Big|-1}\sum_{g\in\hat{\mathcal{G}}^{(d)}_{i}}{\left(g-\hat{\mu}_{\mathcal{G}^{(d)}_{i}}\right)^{2}},$ (21)
@@
$\hat{k}^{(d)}_{i}=\frac{\hat{\mu}_{\mathcal{G}^{(d)}_{i}}^{2}}{\hat{\sigma}^{2}_{\mathcal{G}^{(d)}_{i}}},\quad\hat{\theta}^{(d)}_{i}=\frac{\hat{\sigma}^{2}_{\mathcal{G}^{(d)}_{i}}}{\hat{\mu}_{\mathcal{G}^{(d)}_{i}}}$. (22)
@@
Similarly, the squared Mahalanobis distance from the output embedding to its attributed reference is $\hat{a}_{i}^{(d)}=d^{2}_{M}\left(\bm{\Psi}_{t}^{(d)}(\hat{\mathbf{x}}_{i});\bm{\Psi}_{t}^{(d)}(\mathbf{x}_{i}),\widehat{\widetilde{\bm{\Sigma}}}_{i}^{(d)}\right)$. Consider $Q(k,x)=\Gamma(k,x)/\Gamma(k)$ as the regularized upper incomplete Gamma function (NIST, 2024). Then, the PM score for $\hat{\mathbf{y}}_{i}$ in dimension $d$ is:
$\widehat{\text{PM}}^{(d)}_{i}=Q\!\left(\hat{k}^{(d)}_{i},\frac{\hat{a}_{i}^{(d)}}{\hat{\theta}^{(d)}_{i}}\right),\quad\widehat{\text{PM}}_{i}^{(d)}\in\left[0,1\right],$ (23)